21. You own the following portfolio of stocks. What is the portfolio weight of stock C?

Stock Number of Shares Price per Share
A 110 $22
B 540 $16
C 360 $44
D 190 $36

46.95

56.95

29.95

51.95

39.77
Answer #1
Total value=(Number of shares*Price per share)
A (110*22)=$2420
B (540*16)=$8640
C (360*44)=$15840
D (190*36)=$6840
Total value=$33740

Hence portfolio weight of stock C=(15840/33740)

which is equal to

=46.95%(Approx).